# Robust LLM Alignment via Distributionally Robust Direct Preference Optimization

### v2

@@ A.2 Optimization
Lemma 3 (Beck,, 2014, Theorem 1.24; Linear Approximation Theorem).
Let $f\colon U\to\mathbb{R}$ be a twice continuously differentiable function over an open set $U\subseteq\mathbb{R}^{n}$, and let $x,y\in U$ be such that $[x,y]\subseteq U$. Then there exists $\xi\in[x,y]$ such that
@@
Lemma 4 (Beck,, 2017, Theorem 5.24; First-order characterizations of strong convexity).
@@
Lemma 5 (Beck,, 2017, Theorem 5.25; Existence and uniqueness of a minimizer of closed strongly convex functions).
@@ A.3 Distributionally Robust Optimization Results
[…]
@@
Lemma 6 (Duchi and Namkoong,, 2021, Proposition 1).
Let $D_{f}$ be the $f$-divergence defined in Eq. 12. Then,
@@ A.4 Concentration Results
Lemma 7 (Hoeffding’s inequality (see Boucheron et al.,, 2013, Theorem 2.8)).
Let $X_{1},\dots,X_{n}$ be independent random variables such that $X_{i}$ takes its values in $[a_{i},b_{i}]$ almost surely for all $i\leq n$. Let
@@ Appendix B Proof of WDPO Sample Complexity
Many properties of distributionally robust DPO are derived from those of the non-robust DPO. We hence start with the following proof of policy parameter convergence in the non-robust setting (Proposition 1).
@@ B.1 Proof of Non-robust DPO Policy Parameter Convergence
With $g(\theta)\coloneqq\beta h_{\theta}(s,a^{1},a^{2})$ and the Hessian chain rule for composition with a scalar function (Eq. 14), we have
@@
=$\stackrel{{\scriptstyle(a)}}{{=}}-y\beta\sigma(-\beta h_{\theta}(s,a^{1},a^{2}))\nabla_{\theta}^{2}h_{\theta}(s,a^{1},a^{2})$
@@
In addition, we have that $\nabla_{\theta}^{2}h_{\theta}(s,a^{1},a^{2})=\nabla_{\theta}^{2}\log\pi_{\theta}(a^{1}\mid s)-\nabla_{\theta}^{2}\log\pi_{\theta}(a^{2}\mid s)$. Using the Hessian chain rule (Eq. 14), we have
@@
where $(a)$ is due to Cauchy-Schwarz inequality. $(b)$ is due to the assumptions $\lVert\theta\rVert_{2}\leq B$ and $\max_{s,a}\lVert\psi(s,a)\rVert_{2}\leq 1$. Now this suggests that the input to the function $\sigma(\beta h_{\theta}(s,a^{1},a^{2}))\sigma(-\beta h_{\theta}(s,a^{1},a^{2}))$ is bounded in $[-4\beta B,4\beta B]$. Since $\sigma(x)\sigma(-x)$ is symmetric and strictly decreasing when $x\in[0,\infty)$, we have that
@@
where $\gamma=\frac{\beta^{2}e^{4\beta B}}{(1+e^{4\beta B})^{2}}$ and $X\in\mathbb{R}^{n\times d}$ has the differencing vector $x_{i}\coloneqq\psi(s_{i},a^{1}_{i})-\psi(s_{i},a_{i}^{2})\in\mathbb{R}^{d}$ as its $i$-th row. Thus, if we introduce the error vector $\Delta\coloneqq\theta^{\mathrm{dpo}}_{n}-\theta^{*}$, then by the linear approximation theorem (Lemma 3), there exists $\alpha\in[0,1]$ and $\tilde{\theta}=\alpha\theta^{\mathrm{dpo}}_{n}+(1-\alpha)\theta^{*}$ such that
@@
=$\stackrel{{\scriptstyle(a)}}{{=}}-\beta(y\sigma(\beta h_{\theta}(s,a^{2},a^{1}))-(1-y)\sigma(\beta h_{\theta}(s,a^{1},a^{2})))(\psi(s,a^{1})-\psi(s,a^{2})),$
where $(a)$ is due to $\nabla_{\theta}h_{\theta}(s,a^{1},a^{2})=\psi(s,a^{1})-\psi(s,a^{2})$ calculated in Section B.1. This implies that
@@
Since $X(\Sigma_{\mathcal{D}}+\lambda I)^{-1}X^{\top}$ is symmetric, and clearly $\widetilde{U}\Sigma(\Sigma^{\top}\Sigma/n+\lambda I)^{-1}\Sigma\widetilde{U}^{\top}$ diagonalizes it, the eigenvalue of it takes form $\frac{\sigma_{i}^{2}}{\sigma_{i}^{2}/n+\lambda}$, where $\sigma_{i}$ is the $i$-th singular value of $X$. Hence, all eigenvalues are upper bounded by $n$. Then we must have $\lVert M\rVert_{\mathrm{op}}=\operatorname{\lambda_{\mathrm{max}}}(M)\leq\frac{\beta^{2}}{n}$. Since the components of $V$ are i.i.d. with $\mathbb{E}V_{i}=0$ and $\lvert V_{i}\rvert\leq 1$, the elements are $1$-sub-Gaussian, we can use the Bernstein’s inequality for sub-Gaussian random variables in quadratic form (see Lemma 8). It implies that with probability at least $1-\delta$,
@@ B.2 Proof of WDPO Loss Function Convergence
Recall the strong duality in Lemma 1. The term $\inf_{z\in\mathcal{Z}}[\eta d^{p}(z,z^{\prime})-l(z;\theta)]$ is called the Moreau-Yosida regularization of $-l$ with parameter $1/\eta$. We denote it by $l_{\eta}(z;\theta)$. Now observe that
@@
=$\stackrel{{\scriptstyle(a)}}{{=}}\left\lvert\inf_{\eta\geq 0}\{\eta\rho^{p}-\mathbb{E}_{z\sim\mathsf{P}^{o}}[l_{\eta}(z;\theta)]\}-\inf_{\eta\geq 0}\{\eta\rho^{p}-\mathbb{E}_{z\sim\mathsf{P}_{n}^{o}}[l_{\eta}(z;\theta)]\}\right\rvert$
@@
$-l(z;\theta)=y\log\sigma(x)+(1-y)\log\sigma(-x)\leq 0,\quad y\in\{0,1\}.$
The inequality is because the sigmoid function is strictly bounded between 00 and $1$, i.e., $\sigma\in(0,1)$. This implies that $\log\sigma$ is non-positive. Using this, we have that
@@
Now we prove its lower bound. Recall that in the analysis of non-robust DPO loss, we proved that $\lvert h_{\theta}(s,a^{1},a^{2})\rvert\leq 4B$ (see Eq. 17). Since both $\log$ and $\sigma$ are increasing functions, we have that $\log\sigma(\beta h_{\theta}(s,a^{1},a^{2}))\geq\log\sigma(-4\beta B)$. Now observe that
@@
where the first inequality is because both $\eta$ and metric $d^{p}$ are non-negative. The last inequality is because only one of the $\log\sigma$ term will be activated and the lower bound we recalled above. Denote $K=\lvert\log\sigma(-4\beta B)\rvert$. Since $l_{\eta}$ is a bounded function, by Hoeffding’s inequality for bounded random variable (Lemma 7), we have
@@ B.3 Proof of the Strong Convexity of WDPO Loss
Let $l(z;\theta)$ be the DPO loss function. Assume that Assumption 2 is in place. Then $g(\theta)\coloneqq\mathbb{E}_{z\sim\mathsf{P}}[l(z;\theta)]$ is $\gamma$-strongly convex with respect to norm $\lVert\cdot\rVert_{\Sigma_{\mathsf{P}}}$, where $\Sigma_{\mathsf{P}}=\mathbb{E}_{(s,a^{1},a^{2},y)\sim\mathsf{P}}(\psi(s,a^{1})-\psi(s,a^{2}))(\psi(s,a^{1})-\psi(s,a^{2}))^{\top}$, and $\gamma=\frac{\beta^{2}e^{4\beta B}}{(1+e^{4\beta B})^{2}}$.
@@
In addition, we also proved that (see Eq. 18)
@@
where $\gamma=\frac{\beta^{2}e^{4\beta B}}{(1+e^{4\beta B})^{2}}$. Thus, if we introduce the error vector $\Delta\coloneqq\theta^{\prime}-\theta$, where $\theta,\theta^{\prime}\in\Theta$, then by the linear approximation theorem (Lemma 3), there exists $\alpha\in[0,1]$ and $\tilde{\theta}=\alpha\theta+(1-\alpha)\theta^{\prime}$ such that
@@
where $(a)$ is by Lemma 4. In particular, the equivalence between the inequalities, Eq. 22 and $(a)$, can be found in the proof of Beck, (2017, Theorem 5.24), and the author would like to comment that the proof does not rely on whether $\lVert\cdot\rVert_{\Sigma_{z}}$ is a semi-norm or a norm. Now, by Assumption 2, $\Sigma_{\mathsf{P}}$ is strictly positive definite, hence $\lVert\cdot\rVert_{\Sigma_{\mathsf{P}}}$ is a norm. This implies that $g$ is $\gamma$-strongly convex with respect to $\lVert\cdot\rVert_{\Sigma_{\mathsf{P}}}$. ∎
@@
is $\gamma\lambda$-strongly convex in $\theta$ with respect to (non-weighted) $2$-norm $\lVert\cdot\rVert_{2}$, where $\lambda$ is the regularity condition number defined in Assumption 2, and $\gamma=\frac{\beta^{2}e^{4\beta B}}{(1+e^{4\beta B})^{2}}$.
@@
Let $\alpha\in[0,1]$ and $\theta,\theta^{\prime}\in\Theta$. First, we denote $h(\theta;\mathsf{P})=\mathbb{E}_{z\sim\mathsf{P}}[l(z;\theta)]$ for any $\mathsf{P}$ in the Wasserstein ball. In Lemma 10, we proved that $h$ is $\gamma$-strongly convex in $\theta$ w.r.t. norm $\lVert\cdot\rVert_{\Sigma_{\mathsf{P}}}$. Now observe that
@@
Note that the function $g(\theta)=\mathbb{E}_{z\sim\mathsf{P}}[l(z;\theta)]$ is $\gamma$-strongly convex with respect to $\lVert\cdot\rVert_{\Sigma_{\mathsf{P}}}$ by Lemma 10. We use this fact in $(a)$. The inequality in $(b)$ is due to $\sup_{x}(f(x)+g(x))\leq\sup_{x}f(x)+\sup_{x}g(x)$. The last inequality $(c)$ is because $\operatorname{\lambda_{\mathrm{min}}}(\Sigma_{\mathsf{P}})\geq\lambda$, for all $\mathsf{P}\in\mathcal{P}_{\mathsf{W}}$ by Assumption 2. This implies that $\mathcal{L}^{\mathrm{W}}$ is a $\gamma\lambda$-strongly convex function with respect to $\lVert\cdot\rVert_{2}$. ∎
@@ B.4 Proof of Policy Parameter Convergence of WDPO
By Lemma 9, we have that, with probability at least $1-\delta$,
@@
where the first inequality is because $\theta^{\mathrm{W}}_{n}$ is the minimizer of $\mathcal{L}^{\mathrm{W}}_{n}$. Now by the $\gamma\lambda$-strong convexity of $\mathcal{L}^{\mathrm{W}}$ (see Lemma 11) and Lemma 5.II, we have that
@@ Appendix C Proof of KLDPO Sample Complexity
We state a result from Hu and Hong, (2013) that proves an equivalent condition for the infimum to be achieved at $\lambda^{*}=0$.
Proposition 3 (Hu and Hong,, 2013, Proposition 2).
@@
We include the derivation here for completeness. Previous works in optimization and distributionally robust reinforcement learning have covered the dual problem of distributionally robust optimization with KL uncertainty set (e.g., see Hu and Hong, (2013); Panaganti and Kalathil, (2022); Xu et al., (2023)).
Recall that $f(t)=t\log(t)$ corresponds to the KL divergence. The optimal $t$ for $f^{*}(s)=\sup_{t\geq 0}\{st-t\log(t)\}$ is $\mathrm{exp}\left(s-1\right)$. This implies that the Fenchel conjugate of $f$ is $f^{*}(s)=\mathrm{exp}\left(s-1\right)$. From Lemma 6, we get
@@
This implies that it suffices to seek the infimum in $[0,\overline{\lambda}_{\theta}]$. Hence, we have
@@
Now from Proposition 3, the condition $\log\kappa_{u}+\rho\geq 0$ is problem-dependent due to the diameter $\rho$, which is a design choice. Note that when $\kappa_{u}$ is close to zero, the condition $\log\kappa_{u}+\rho\geq 0$ is almost never true for a reasonable $\rho$. Hence, we ignore the case where $\lambda^{*}=0$. By Assumption 3, without loss of generality, we have that $\lambda^{*}\in[\underline{\lambda},\overline{\lambda}]$, where $\underline{\lambda}$ is some problem-specific constant. Then we have the result. In the literature of distributionally robust reinforcement learning, similar arguments can be found in Zhou et al., (2021); Panaganti and Kalathil, (2022). ∎
@@
Fix any $\theta\in\Theta$ and $\rho>0$. Let $\delta\in(0,1)$. Assume Assumption 3 is in place. With probability $1-\delta$, we have that
@@
=$\stackrel{{\scriptstyle(a)}}{{=}}\bigg{\lvert}\inf_{\lambda\in[\underline{\lambda},\overline{\lambda}]}\bigg{\{}\lambda\rho+\lambda\log\bigg{(}\mathbb{E}_{z\sim\mathsf{P}^{o}}\bigg{[}\mathrm{exp}\left(\frac{l(z;\theta)}{\lambda}\right)\bigg{]}\bigg{)}\bigg{\}}$
@@
=$\stackrel{{\scriptstyle(c)}}{{=}}\sup_{\lambda\in[\underline{\lambda},\overline{\lambda}]}\lambda\left\lvert\log\bigg{(}\frac{\mathbb{E}_{z\sim\mathsf{P}^{o}_{n}}[\mathrm{exp}\left(l(z;\theta)\right)/\lambda]}{\mathbb{E}_{z\sim\mathsf{P}^{o}}[\mathrm{exp}\left(l(z;\theta)\right)/\lambda]}\bigg{)}\right\rvert$
@@
where $(a)$ is by Lemma 12. $(b)$ is because $\lvert\inf_{x}f(x)-\inf_{x}g(x)\rvert\leq\sup_{x}\lvert f(x)-g(x)\rvert$. $(c)$ is by Assumption 3. $(d)$ is due to $\lvert\log(1+x)\rvert\leq\lvert x\rvert,\forall x\geq 0$. $(e)$ is due to the fact that the loss function $l$ is non-negative, i.e., $l\geq 0$. Now by applying Hoeffding’s inequality (Lemma 7), we have
@@
We prove a strong convexity result similar to Lemma 11 for KLDPO loss function.
@@
is $\gamma\lambda$-strongly convex in $\theta$ with respect to (non-weighted) $2$-norm $\lVert\cdot\rVert_{2}$, where $\lambda$ is the regularity condition number defined in Assumption 2, and $\gamma=\frac{\beta^{2}e^{4\beta B}}{(1+e^{4\beta B})^{2}}$.
@@
Let $\alpha\in[0,1]$ and $\theta,\theta^{\prime}\in\Theta$. First, we denote $h(\theta;\mathsf{P})=\mathbb{E}_{z\sim\mathsf{P}}[l(z;\theta)]$ for any $\mathsf{P}$ in the KL ball. In Lemma 10, we proved that $h$ is $\gamma$-strongly convex in $\theta$ w.r.t. norm $\lVert\cdot\rVert_{\Sigma_{\mathsf{P}}}$. Now observe that

### v3

@@ A.2 Optimization
Lemma 3 (Beck, 2014, Theorem 1.24; Linear Approximation Theorem).
Let $f\colon U\to\mathbb{R}$ be a twice continuously differentiable function over an open set $U\subseteq\mathbb{R}^{n}$, and let $x,y\in U$ be such that $x,y]\subseteq U$. Then there exists $\xi\in[x,y]$ such that
@@
Lemma 4 (Beck, 2017, Theorem 5.24; First-order characterizations of strong convexity).
@@
Lemma 5 (Beck, 2017, Theorem 5.25; Existence and uniqueness of a minimizer of closed strongly convex functions).
@@ A.3 Distributionally Robust Optimization Results
$D_{f}(\mathsf{P}\;\|\;\mathsf{P}^{o})=\int_{\mathcal{X}}f\bigg(\frac{d\mathsf{P}}{d\mathsf{P}^{o}}\bigg)d\mathsf{P}^{o},$ (12)
@@
Lemma 6 (Duchi and Namkoong, 2021, Proposition 1).
Let $D_{f}$ be the $f$-divergence defined in Eq.˜12. Then,
@@ A.4 Concentration Results
Lemma 7 (Hoeffding’s inequality (see Boucheron et al., 2013, Theorem 2.8)).
Let $X_{1},\dots,X_{n}$ be independent random variables such that $X_{i}$ takes its values in $a_{i},b_{i}]$ almost surely for all $i\leq n$. Let
@@ Appendix B Proof of WDPO Sample Complexity
Many properties of distributionally robust DPO are derived from those of the non-robust DPO. We hence start with the following proof of policy parameter convergence in the non-robust setting (Proposition˜1).
@@ B.1 Proof of Non-robust DPO Policy Parameter Convergence
With $g(\theta)\coloneqq\beta h_{\theta}(s,a^{1},a^{2})$ and the Hessian chain rule for composition with a scalar function (Eq.˜14), we have
@@
$\stackrel{{\scriptstyle(a)}}{{=}}-y\beta\sigma(-\beta h_{\theta}(s,a^{1},a^{2}))\nabla_{\theta}^{2}h_{\theta}(s,a^{1},a^{2})$
@@
In addition, we have that $\nabla_{\theta}^{2}h_{\theta}(s,a^{1},a^{2})=\nabla_{\theta}^{2}\log\pi_{\theta}(a^{1}\mid s)-\nabla_{\theta}^{2}\log\pi_{\theta}(a^{2}\mid s)$. Using the Hessian chain rule (Eq.˜14), we have
@@
where $(a)$ is due to Cauchy-Schwarz inequality. $(b)$ is due to the assumptions $\lVert\theta\rVert_{2}\leq B$ and $\max_{s,a}\lVert\psi(s,a)\rVert_{2}\leq 1$. Now this suggests that the input to the function $\sigma(\beta h_{\theta}(s,a^{1},a^{2}))\sigma(-\beta h_{\theta}(s,a^{1},a^{2}))$ is bounded in $4\beta B,4\beta B]$. Since $\sigma(x)\sigma(-x)$ is symmetric and strictly decreasing when $x\in[0,\infty)$, we have that
@@
where $\gamma=\frac{\beta^{2}e^{4\beta B}}{(1+e^{4\beta B})^{2}}$ and $X\in\mathbb{R}^{n\times d}$ has the differencing vector $x_{i}\coloneqq\psi(s_{i},a^{1}_{i})-\psi(s_{i},a_{i}^{2})\in\mathbb{R}^{d}$ as its $i$-th row. Thus, if we introduce the error vector $\Delta\coloneqq\theta^{\mathrm{dpo}}_{n}-\theta^{*}$, then by the linear approximation theorem (Lemma˜3), there exists $\alpha\in[0,1]$ and $\tilde{\theta}=\alpha\theta^{\mathrm{dpo}}_{n}+(1-\alpha)\theta^{*}$ such that
@@
$\stackrel{{\scriptstyle(a)}}{{=}}-\beta(y\sigma(\beta h_{\theta}(s,a^{2},a^{1}))-(1-y)\sigma(\beta h_{\theta}(s,a^{1},a^{2})))(\psi(s,a^{1})-\psi(s,a^{2})),$
where $(a)$ is due to $\nabla_{\theta}h_{\theta}(s,a^{1},a^{2})=\psi(s,a^{1})-\psi(s,a^{2})$ calculated in Section˜B.1. This implies that
@@
Since $X(\Sigma_{\mathcal{D}}+\lambda I)^{-1}X^{\top}$ is symmetric, and clearly $\widetilde{U}\Sigma(\Sigma^{\top}\Sigma/n+\lambda I)^{-1}\Sigma\widetilde{U}^{\top}$ diagonalizes it, the eigenvalue of it takes form $\frac{\sigma_{i}^{2}}{\sigma_{i}^{2}/n+\lambda}$, where $\sigma_{i}$ is the $i$-th singular value of $X$. Hence, all eigenvalues are upper bounded by $n$. Then we must have $\lVert M\rVert_{\mathrm{op}}=\operatorname{\lambda_{\mathrm{max}}}(M)\leq\frac{\beta^{2}}{n}$. Since the components of $V$ are i.i.d. with $\mathbb{E}V_{i}=0$ and $\lvert V_{i}\rvert\leq 1$, the elements are $1$-sub-Gaussian, we can use the Bernstein’s inequality for sub-Gaussian random variables in quadratic form (see Lemma˜8). It implies that with probability at least $1-\delta$,
@@ B.2 Proof of WDPO Loss Function Convergence
Recall the strong duality in Lemma˜1. The term $\inf_{z\in\mathcal{Z}}[\eta d^{p}(z,z^{\prime})-l(z;\theta)]$ is called the Moreau-Yosida regularization of l-l with parameter $1/\eta$. We denote it by $l_{\eta}(z;\theta)$. Now observe that
@@
$\stackrel{{\scriptstyle(a)}}{{=}}\left\lvert\inf_{\eta\geq 0}\{\eta\rho^{p}-\mathbb{E}_{z\sim\mathsf{P}^{o}}[l_{\eta}(z;\theta)]\}-\inf_{\eta\geq 0}\{\eta\rho^{p}-\mathbb{E}_{z\sim\mathsf{P}_{n}^{o}}[l_{\eta}(z;\theta)]\}\right\rvert$
@@
$l(z;\theta)=y\log\sigma(x)+(1-y)\log\sigma(-x)\leq 0,\quad y\in\{0,1\}.$
The inequality is because the sigmoid function is strictly bounded between 0 and $1$, i.e., $\sigma\in(0,1)$. This implies that $\log\sigma$ is non-positive. Using this, we have that
@@
Now we prove its lower bound. Recall that in the analysis of non-robust DPO loss, we proved that $\lvert h_{\theta}(s,a^{1},a^{2})\rvert\leq 4B$ (see Eq.˜17). Since both $\log$ and $\sigma$ are increasing functions, we have that $\log\sigma(\beta h_{\theta}(s,a^{1},a^{2}))\geq\log\sigma(-4\beta B)$. Now observe that
@@
where the first inequality is because both $\eta$ and metric $d^{p}$ are non-negative. The last inequality is because only one of the $\log\sigma$ term will be activated and the lower bound we recalled above. Denote $K=\lvert\log\sigma(-4\beta B)\rvert$. Since $l_{\eta}$ is a bounded function, by Hoeffding’s inequality for bounded random variable (Lemma˜7), we have
@@ B.3 Proof of the Strong Convexity of WDPO Loss
Let $l(z;\theta)$ be the DPO loss function. Assume that Assumption˜2 is in place. Then $g(\theta)\coloneqq\mathbb{E}_{z\sim\mathsf{P}}[l(z;\theta)]$ is $\gamma$-strongly convex with respect to norm $\lVert\cdot\rVert_{\Sigma_{\mathsf{P}}}$, where $\Sigma_{\mathsf{P}}=\mathbb{E}_{(s,a^{1},a^{2},y)\sim\mathsf{P}}(\psi(s,a^{1})-\psi(s,a^{2}))(\psi(s,a^{1})-\psi(s,a^{2}))^{\top}$, and $\gamma=\frac{\beta^{2}e^{4\beta B}}{(1+e^{4\beta B})^{2}}$.
@@
In addition, we also proved that (see Eq.˜18)
@@
where $\gamma=\frac{\beta^{2}e^{4\beta B}}{(1+e^{4\beta B})^{2}}$. Thus, if we introduce the error vector $\Delta\coloneqq\theta^{\prime}-\theta$, where $\theta,\theta^{\prime}\in\Theta$, then by the linear approximation theorem (Lemma˜3), there exists $\alpha\in[0,1]$ and $\tilde{\theta}=\alpha\theta+(1-\alpha)\theta^{\prime}$ such that
@@
where $(a)$ is by Lemma˜4. In particular, the equivalence between the inequalities, Eq.˜22 and $(a)$, can be found in the proof of Beck (2017, Theorem 5.24), and the author would like to comment that the proof does not rely on whether $\lVert\cdot\rVert_{\Sigma_{z}}$ is a semi-norm or a norm. Now, by Assumption˜2, $\Sigma_{\mathsf{P}}$ is strictly positive definite, hence $\lVert\cdot\rVert_{\Sigma_{\mathsf{P}}}$ is a norm. This implies that $g$ is $\gamma$-strongly convex with respect to $\lVert\cdot\rVert_{\Sigma_{\mathsf{P}}}$. ∎
@@
is $\gamma\lambda$-strongly convex in $\theta$ with respect to (non-weighted) $2$-norm $\lVert\cdot\rVert_{2}$, where $\lambda$ is the regularity condition number defined in Assumption˜2, and $\gamma=\frac{\beta^{2}e^{4\beta B}}{(1+e^{4\beta B})^{2}}$.
@@
Let $\alpha\in[0,1]$ and $\theta,\theta^{\prime}\in\Theta$. First, we denote $h(\theta;\mathsf{P})=\mathbb{E}_{z\sim\mathsf{P}}[l(z;\theta)]$ for any $\mathsf{P}$ in the Wasserstein ball. In Lemma˜10, we proved that $h$ is $\gamma$-strongly convex in $\theta$ w.r.t. norm $\lVert\cdot\rVert_{\Sigma_{\mathsf{P}}}$. Now observe that
@@
Note that the function $g(\theta)=\mathbb{E}_{z\sim\mathsf{P}}[l(z;\theta)]$ is $\gamma$-strongly convex with respect to $\lVert\cdot\rVert_{\Sigma_{\mathsf{P}}}$ by Lemma˜10. We use this fact in $(a)$. The inequality in $(b)$ is due to $\sup_{x}(f(x)+g(x))\leq\sup_{x}f(x)+\sup_{x}g(x)$. The last inequality $(c)$ is because $\operatorname{\lambda_{\mathrm{min}}}(\Sigma_{\mathsf{P}})\geq\lambda$, for all $\mathsf{P}\in\mathcal{P}_{\mathsf{W}}$ by Assumption˜2. This implies that $\mathcal{L}^{\mathrm{W}}$ is a $\gamma\lambda$-strongly convex function with respect to $\lVert\cdot\rVert_{2}$. ∎
@@ B.4 Proof of Policy Parameter Convergence of WDPO
By Lemma˜9, we have that, with probability at least $1-\delta$,
@@
where the first inequality is because $\theta^{\mathrm{W}}_{n}$ is the minimizer of $\mathcal{L}^{\mathrm{W}}_{n}$. Now by the $\gamma\lambda$-strong convexity of $\mathcal{L}^{\mathrm{W}}$ (see Lemma˜11) and Lemma˜5.II, we have that
@@ Appendix C Proof of KLDPO Sample Complexity
We state a result from Hu and Hong (2013) that proves an equivalent condition for the infimum to be achieved at $\lambda^{*}=0$.
Proposition 3 (Hu and Hong, 2013, Proposition 2).
@@
We include the derivation here for completeness. Previous works in optimization and distributionally robust reinforcement learning have covered the dual problem of distributionally robust optimization with KL uncertainty set (e.g., see Hu and Hong (2013); Panaganti and Kalathil (2022); Xu et al. (2023)).
Recall that $f(t)=t\log(t)$ corresponds to the KL divergence. The optimal $t$ for $f^{*}(s)=\sup_{t\geq 0}\{st-t\log(t)\}$ is $\mathrm{exp}\left(s-1\right)$. This implies that the Fenchel conjugate of $f$ is $f^{*}(s)=\mathrm{exp}\left(s-1\right)$. From Lemma˜6, we get
@@
This implies that it suffices to seek the infimum in $0,\overline{\lambda}_{\theta}]$. Hence, we have
@@
Now from Proposition˜3, the condition $\log\kappa_{u}+\rho\geq 0$ is problem-dependent due to the diameter $\rho$, which is a design choice. Note that when $\kappa_{u}$ is close to zero, the condition $\log\kappa_{u}+\rho\geq 0$ is almost never true for a reasonable $\rho$. Hence, we ignore the case where $\lambda^{*}=0$. By Assumption˜3, without loss of generality, we have that $\lambda^{*}\in[\underline{\lambda},\overline{\lambda}]$, where $\underline{\lambda}$ is some problem-specific constant. Then we have the result. In the literature of distributionally robust reinforcement learning, similar arguments can be found in Zhou et al. (2021); Panaganti and Kalathil (2022). ∎
@@
Fix any $\theta\in\Theta$ and $\rho>0$. Let $\delta\in(0,1)$. Assume Assumption˜3 is in place. With probability $1-\delta$, we have that
@@
$\stackrel{{\scriptstyle(a)}}{{=}}\bigg\lvert\inf_{\lambda\in[\underline{\lambda},\overline{\lambda}]}\bigg\{\lambda\rho+\lambda\log\bigg(\mathbb{E}_{z\sim\mathsf{P}^{o}}\bigg[\mathrm{exp}\left(\frac{l(z;\theta)}{\lambda}\right)\bigg]\bigg)\bigg\}$
@@
$\stackrel{{\scriptstyle(c)}}{{=}}\sup_{\lambda\in[\underline{\lambda},\overline{\lambda}]}\lambda\left\lvert\log\bigg(\frac{\mathbb{E}_{z\sim\mathsf{P}^{o}_{n}}[\mathrm{exp}\left(l(z;\theta)\right)/\lambda]}{\mathbb{E}_{z\sim\mathsf{P}^{o}}[\mathrm{exp}\left(l(z;\theta)\right)/\lambda]}\bigg)\right\rvert$
@@
where $(a)$ is by Lemma˜12. $(b)$ is because $\lvert\inf_{x}f(x)-\inf_{x}g(x)\rvert\leq\sup_{x}\lvert f(x)-g(x)\rvert$. $(c)$ is by Assumption˜3. $(d)$ is due to $\lvert\log(1+x)\rvert\leq\lvert x\rvert,\forall x\geq 0$. $(e)$ is due to the fact that the loss function $l$ is non-negative, i.e., $l\geq 0$. Now by applying Hoeffding’s inequality (Lemma˜7), we have
@@
We prove a strong convexity result similar to Lemma˜11 for KLDPO loss function.
@@
is $\gamma\lambda$-strongly convex in $\theta$ with respect to (non-weighted) $2$-norm $\lVert\cdot\rVert_{2}$, where $\lambda$ is the regularity condition number defined in Assumption˜2, and $\gamma=\frac{\beta^{2}e^{4\beta B}}{(1+e^{4\beta B})^{2}}$.
@@
Let $\alpha\in[0,1]$ and $\theta,\theta^{\prime}\in\Theta$. First, we denote $h(\theta;\mathsf{P})=\mathbb{E}_{z\sim\mathsf{P}}[l(z;\theta)]$ for any $\mathsf{P}$ in the KL ball. In Lemma˜10, we proved that $h$ is $\gamma$-strongly convex in $\theta$ w.r.t. norm $\lVert\cdot\rVert_{\Sigma_{\mathsf{P}}}$. Now observe that